# UPRE: Zero-Shot Domain Adaptation for Object Detection via Unified Prompt and Representation Enhancement

### v1

@@ 1 Introduction
Domain adaptation has gained widespread attention in recent years [21, 4, 37, 48, 40, 39, 33, 30]. However, obtaining image priors, even unlabeled, is not always feasible, thereby limiting their applicability in real-world scenarios. Zero-shot domain adaptation (ZSDA) [38, 58, 50, 13, 60, 2] aims to address this constraint by enabling adaptation to target domains without prior image exposure. The development of vision-language models (VLMs) [32, 31, 1, 24, 54, 7] has markedly advanced ZSDA, by utilizing textual prompts to describe unseen domains through their inherent zero-shot capabilities [13, 60]. Despite these advancements, our analysis reveals two primary limitations in existing VLM-based approaches: 1) domain bias, where distribution shifts between source and target domains introduces task-agnostic noise, impairing model performance; 2) detection bias, where models such as CLIP [42] emphasize global image representations but overlook instance-level details crucial for precise object localization. This deficiency stems from manually constructed prompts, which inadequately capture the contextual attributes of foreground and background objects.
@@ 4.4 Ablation Study
Methods Daytime Foggy Night Clear Night Rainy Dusk Rainy
$\mathcal{L}_{ce}$ 37.0 37.3 16.5 30.8
$\mathcal{L}_{{{bg}}}$ only 36.6 35.7 15.5 29.6
$\mathcal{L}_{c}$ only 37.6 37.8 17.1 31.4
$\mathcal{L}_{{{bg}}}$ + $\mathcal{L}_{c}$ 38.7 38.9 17.8 32.8
@@ B Prompt Templates
For a fair comparison, we adopt the same prompts used in CLIP [42]. Following previous works [47, 28], we extend the category names and domain characteristics into sentences using multi-view prompts at both the image and instance levels. Specifically, we use ”A photo taken in a $[domain]$.” as the image-level prompt input for the text encoder. Then, we define ”A $[domain]$ photo of a $[class]$.” as the instance-level positive prompt and ”A $[domain]$ photo of an $[unknown\;\,class]$.” as the instance-level negative prompt. For the three cross-domain scenarios, we employ 90 $[domain]$-specific prompt templates, as illustrated in Figure 1. We apply L2 normalization to obtain the final multi-view prompt representations.

### v2

@@ 1 Introduction
Domain adaptation has gained widespread attention in recent years [23, 4, 41, 52, 44, 43, 36, 33]. However, obtaining image priors, even unlabeled, is not always feasible, thereby limiting their applicability in real-world scenarios. Zero-shot domain adaptation (ZSDA) [42, 62, 54, 14, 65, 2] aims to address this constraint by enabling adaptation to target domains without prior image exposure. The development of vision-language models (VLMs) [35, 34, 1, 27, 9, 58, 7, 38] has markedly advanced ZSDA, by utilizing textual prompts to describe unseen domains through their inherent zero-shot capabilities [14, 65]. Despite these advancements, our analysis reveals two primary limitations in existing VLM-based object detection approaches: 1) domain bias, where distribution shifts between source and target domains introduces task-agnostic noise, impairing model performance; 2) detection bias, where models such as CLIP [46] emphasize global image representations but overlook instance-level details crucial for precise object localization. This deficiency stems from manually constructed prompts, which inadequately capture the contextual attributes of foreground and background objects.
@@ 4.3 Main Results
Refer to caption Refer to caption
Figure 5: The t-SNE visualization of image embeddings across five domains using identical images and experimental settings. CLIP (Left) represents coarse generalization capabilities, whereas UPRE (Right) exhibits superior adaptation to each target domain.
The visualization of the domain embeddings. As depicted in Fig. 5 (left), although CLIP possesses generalization abilities, it lacks discrimination in embeddings for different weather conditions, such as the entanglement of Night Clear and Night Rainy embeddings. In Fig. 5 (right), UPRE enhances CLIP’s performance and achieves superior domain adaptation capability.
@@ A Additional Ablation Studies
Methods Daytime Foggy Night Clear Night Rainy Dusk Rainy
$\mathcal{L}_{ce}$ 37.0 37.3 16.5 30.8
$\mathcal{L}_{{{bg}}}$ only 36.6 35.7 15.5 29.6
$\mathcal{L}_{c}$ only 37.6 37.8 17.1 31.4
$\mathcal{L}_{{{bg}}}$ + $\mathcal{L}_{c}$ 38.7 38.9 17.8 32.8
Method Prompt Category Label of Daytime Night Night Dusk
Design Space Negatives Foggy Clear Rainy Rainy
DetPro Negative $\mathcal{C}\cup{\mathcal{C}_{bg}}$ 0 or 1 38.1 38.4 19.0 32.8
DetPro Shared $\mathcal{C}$ $\frac{1}{|\mathcal{C}|}$ 38.5 37.9 18.8 33.3
PNS (Ours) Negative $\mathcal{C}\cup{\mathcal{C}_{bg}}$ $\frac{1}{|\mathcal{C}\cup{\mathcal{C}_{bg}}|}$ 40.0 41.5 19.8 34.5
Table 15: DetPro vs. PNS: comparative analysis and ablation study.
Method Category Inference Computation Total Daytime Daytime Night Night Dusk
Time Cost Parameters Clear Foggy Clear Rainy Rainy
Faster-RCNN Traditional 67ms 183G 52.7M 48.1 32.0 34.4 12.4 26.0
OA-DG 69ms 183G 52.7M 55.8 38.3 38.0 16.8 33.9
DAI-Net* 124ms 297G 78.1M 54.4 36.7 41.0 18.9 33.0
UFR - - - 58.6 39.6 40.8 19.2 33.2
CLIP-GAP VLM-based 98ms 526G 131.4M 51.3 38.5 36.9 18.7 32.3
PODA* 72ms 185G 129.3M 51.8 39.2 38.7 19.0 33.4
PDD 101ms 531G 134.3M 53.6 39.1 38.5 19.2 33.7
UPRE(Ours) VLM-based 111ms 528G 129.8M 53.9 40.0 41.5 19.8 34.5
Table 16: The comparison of efficiency and effectiveness. All test settings are same. CLIP-GAP, PDD and UPRE are all based on the Detectron2 framework, using CLIP’s ResNet101 backbone. UFR’s - denotes code and model remain unavailable. PODA only use visual encoder during inference.
Comparison with DetPro. In Table 15, we present ablation studies that highlight differences between PNS and DetPro. As shown in line 1, DetPro struggles with training negative prompt, because it only learns a prompt embedding that draws all negative proposals close with a simple label of 0 or 1 (Detpro’s Eq.(8)). Therefore, as depicted in line 2, DetPro opts to train the shared prompt by forcing negative proposals to be equally unlike any foreground classes in category space $\mathcal{C}$ (Detpro’s Eq.(5)). However, selecting from foreground classes with a probability of $\frac{1}{|\mathcal{C}|}$ overlooks the utilization of the background category.
In contrast to DetPro, PNS trains negative prompt in category space $\mathcal{C}\cup{\mathcal{C}_{bg}}$, using $\frac{1}{|\mathcal{C}\cup{\mathcal{C}_{bg}}|}$ for labeling negative proposals in Eq.(12). This approach allows negative proposals to be equally unlike any classes in Eq.(11), which aids the negative prompt in learning diverse context, as negative proposals variably encompass either pure backgrounds or parts of objects. Contrasted with DetPro’s shared and negative prompts, our method excels in performance across all target domains, achieving notable improvement by 3.6% and 3.1% in Night Clear, respectively. DetPro insufficiently explores negative prompt, while PNS effectively trains the negative prompt. PNS is not a direct application of DetPro; it innovatively explores proposal separation in training negative prompt. Moreover, proposal separation is a commonly used trick in detection tasks.
Efficiency analysis. Applying advanced large models [64, 46, 37] to ZSDA is promising but inevitably faces higher computational and memory costs. As illustrated in Table 16, UPRE outshines VLM-based methods, notably PDD, with 4.5M fewer parameters and a reduction of 3 GFLOPs in computational cost. Although UPRE is 10ms slower than PDD, this latency is acceptable given its superior performance. To ensure efficiency, we only leverage the MDP module during inference, which uses less than 0.1M parameters.
Evidence of improvement attributed to domain adaptation. To assess whether improvements stem from better domain adaptation [20, 25] or just a stronger detector, we compare different methods on the source domain to directly reflect detector performance. As shown in the Daytime clear column of Table 16, UPRE performs moderately but shows significant improvement in target domains.
@@ B Prompt Templates
For a fair comparison, we adopt the same prompts used in CLIP [46]. Following previous works [51, 31], we extend the category names and domain characteristics into sentences using multi-view prompts at both the image and instance levels. Specifically, we use "A photo taken in a $[domain]$." as the image-level prompt input for the text encoder. Then, we define "A $[domain]$ photo of a $[class]$." as the instance-level positive prompt and "A $[domain]$ photo of an $[unknown\;\,class]$." as the instance-level negative prompt. For the three cross-domain scenarios, we employ 90 $[domain]$-specific prompt templates, as illustrated in Figure 1. We apply L2 normalization to obtain the final multi-view prompt representations.